# Junchen Fu1 Xuri Ge2 Corresponding author. Kaiwen Zheng1 Ioannis Arapakis3 Xin Xin2 Joemon M. Jose1

### v2

@@ Abstract
Popular Micro-videos, dominant on platforms like TikTok and YouTube, hold significant commercial value. The rise of high-quality AI-generated content has spurred interest in AI-driven micro-video creation. However, despite the advanced capabilities of large language models (LLMs) like ChatGPT and DeepSeek in text generation and reasoning, their potential to assist the creation of popular micro-videos remains largely unexplored.
In this paper, we conduct an empirical study on LLM-assisted popular micro-video generation (LLMPopcorn [^1] We selected popcorn as the icon for this paper because it symbolizes leisure and entertainment. This aligns with this study on leveraging LLMs as assistants for generating popular micro-videos, which are often consumed during leisure time.). Specifically, we investigate the following research questions: (i) How can LLMs be effectively utilized to assist popular micro-video generation? (ii) To what extent can prompt-based enhancements optimize the LLM-generated content for higher popularity? (iii) How well do various LLMs and video generators perform in the popular micro-video generation task? By exploring these questions, we show that advanced LLMs like DeepSeek-V3 enable micro-video generation to achieve popularity comparable to human-created content. Prompt enhancements further boost popularity, and benchmarking highlights DeepSeek-V3 and DeepSeek-R1 among LLMs, while LTX-Video and HunyuanVideo lead in video generation. This pioneering work advances AI-assisted micro-video creation, uncovering new research opportunities. We will release the code and datasets to support future studies.
[Uncaptioned image]     LLMPopcorn: An Empirical Study of LLMs as Assistants for Popular Micro-video Generation
@@ 1 Introduction
Micro-videos (or short videos) have emerged as a fundamental element of the digital economy, representing a multi-billion-dollar industry Guo et al. (2024). They have become an integral part of daily life for people worldwide, providing substantial commercial value for social media platforms and content creators. Popular content creators can receive significant revenue through their content Ran et al. (2022), which underscores the ever-growing influence of micro-videos in modern society.
@@
Despite their widespread popularity and financial impact, producing popular micro-video content remains a costly and labor-intensive process. Professional filming, scripting, and editing require significant time and resources, which not all creators or businesses can afford. Driven by these challenges, we explore simplifying the micro-video creation process using AI solutions, with a particular focus on generating popular micro-videos.
Figure 1: An overview of the LLMPopcorn pipeline. The numbered steps indicate the sequential implementation of the pipeline. Given a user query as input, the pipeline generates video titles and micro-videos, followed by an evaluation process.
@@
On one hand, the rapid advancement of Large Language Models (LLMs) like ChatGPT has unlocked new possibilities for content generation, demonstrating strong capabilities in tasks such as document summarization Jin et al. (2024), programming Cai et al. (2023); Ouyang et al. (2023), mathematical reasoning Shao et al. (2024); Zhang et al. (2024a) and recommendation Yuan et al. (2023); Fu et al. (2024a, b, c). On the other hand, video generation models powered by diffusion techniques and neural rendering are transforming creative content production, enabling high-quality synthesis for applications in film production Liu et al. (2024b) and interactive media Zhang et al. (2024b). Significant breakthroughs in LLMs and video generation models have prompted many studies Zhou et al. (2024) to integrate LLM to facilitate the capability of automatic video generation. For instance, VideoLLMs like VideoPoet Kondratyuk et al. (2023) and GPT4Video Wang et al. (2024) leverage LLMs and video generation models for multimodal video understanding and generation.
Although LLMs can assist in video generation, existing studies have neglected the ability of different LLMs to assist in the generation of “popular” micro-videos. The main reason for this is that existing studies focus on the quality of video generation, such as high resolution, long duration, etc., but neglect that popularity of video generation does not only depend on these. Recently, many studies Cho et al. (2024); Zhong et al. (2024); Lu et al. (2024) have started to focus on and propose the latest micro-video popularity prediction models, which comprehensively integrate micro-video popularity factors. For instance, the Microlens dataset (Ni et al., 2023), which includes raw videos and popularity metrics, along with pre-trained short video popularity prediction models (Zhong et al., 2024), provides a strong basis for video popularity prediction. However, how mainstream LLMs can be utilised to facilitate popular micro-video generation, especially in effectively measuring the high-popularity of generated videos, remains largely unexamined by the research community.
@@
Driven by the above insights, we explore the research question of whether LLMs can assist in the popular micro-video creation. To address this research question systematically, we propose and investigate three following Research Questions (RQs):
@@
  • RQ1: How can LLMs be effectively utilized to assist popular micro-video generation? We address this by developing LLMPopcorn, a comprehensive popular micro-video creation pipeline that leverages LLMs for generating prompts tailored to popular video production. To introduce video popularity assessment in the pipeline, we employ a state-of-the-art offline popularity evaluator to quantify the popularity of generated videos. We compare three mainstream LLMs in our LLMPopcorn pipeline and evaluate their performance alongside human-created videos.
  • RQ2: To what extent can prompt-based enhancements optimize the content generated by LLMs for achieving greater popularity? To address the question, we propose and explore a new Prompt Enhancement that leverages Retrieval-Augmented Generation (RAG) and chain-of-thought (CoT) prompting. This approach is inspired by the video creation process typical of human content creators.
@@
  • RQ3: How do various LLMs and video generators perform on popular micro-video generation? Five state-of-the-art LLMs and three video generation models under the LLMPopcorn pipeline fully compare the performance of different models.
@@
Figure 2: An overview of the Prompt Enhancement (PE) process. PE enables the LLM reviewing relevant micro-videos from the database and engaging in chain-of-thought reasoning.
@@
Anime Summarize the top 5 trending anime this season with a brief analysis.
Create a video exploring how anime portrays futuristic cities.
Generate a video analyzing the symbolism in anime opening themes.
Create a video showcasing Goku’s most iconic transformations in Dragon Ball.
Animate a battle between Levi Ackerman and a Titan in Attack on Titan.
Make a montage of Naruto’s Rasengan evolution throughout the series.
Delicacy Create a video about the history and evolution of a popular street food.
Showcase 5 unique ways desserts are prepared across different cultures.
Compare how breakfast is eaten in various countries.
Generate a video showcasing the preparation of wagyu beef steak.
Create a video about assembling a colorful bento box.
Show the step-by-step process of making matcha tiramisu.
Daily Sharing Create a video showcasing life hacks that simplify daily routines.
Generate a video on organizing spaces for maximum efficiency.
Explore tips for creating a morning routine that boosts productivity.
Generate a video showing the most creative DIY desk setups for students.
Create a video of organizing a messy kitchen cabinet with clever hacks.
Highlight a daily skincare routine featuring Korean beauty products.
Film and TV Generate a video ranking iconic moments in cinematic history.
Explore how directors use symbolism in lighting and shadows.
Create a video comparing film adaptations of popular novels.
Generate a video ranking the top five Marvel Cinematic Universe villains.
Create a video comparing the different Batmen from movie history.
Highlight the most iconic scenes in The Lord of the Rings trilogy.
Game Create a video analyzing how games portray complex moral choices.
Generate a list of the most innovative game mechanics in recent years.
Discuss how multiplayer games foster social interactions.
Generate a video showing the top 10 sniper rifles in Call of Duty history.
Create a video ranking the most popular Minecraft mods.
Highlight the funniest glitches in Skyrim.
Table 1: User prompt examples for abstract and concrete video generation tasks.
2 The Proposed LLMPopcorn
2.1 Problem Formulation
We consider a problem in which a user provides an input prompt $x\in\mathcal{X}$. An LLM processes this prompt and generates both a video title and a video prompt, which we denote by $f(x)=(t,p)$, where $t$ belongs to the set of titles $\mathcal{T}$ and $p$ belongs to the set of video prompts $\mathcal{P}$. Next, a pre-trained video generation model $g$ uses the generated prompt to produce an AI-created micro-video $v\in\mathcal{V}$, that is, $v=g(p)$. An offline video popularity predictor $h$ then assigns a popularity score $s\in\mathbb{R}$ to the generated video, so that $s=h(v,t)$. Overall, the entire pipeline can be represented as the composite function:
$F(x)=h\bigl{(}g(f(x))\bigr{)}.$
The objective is to optimize the LLM function $f$ (with $g$ and $h$ fixed) to maximize the expected popularity score of the generated videos. Formally, we wish to achieve:
$\max_{f\in\mathcal{F}}\;\mathbb{E}_{x\sim\mathcal{X}}\left[h\bigl{(}g(f(x))\bigr{)}\right],$
where $\mathcal{F}$ denotes the set of all possible configurations of the LLM. This formulation captures the aim of leveraging LLMs to enhance the creation of popular AI-generated micro-videos.
2.2 Pipeline
Based on the problem formulation, we propose a new LLMPopcorn pipeline in Figure 1 that begins by constructing user prompts, categorized into concrete and abstract prompts, which are constructed by rules into inputs for an LLM. We demonstrate the basic prompt template in Appendix Figure 9. The LLM serves as a video generation assistant, producing video titles and video generation prompts based on these inputs. These video generation prompts are then fed into a video generator to create corresponding videos. The generated videos are combined with their titles to form final short videos, which are subsequently evaluated using a pre-trained video popularity prediction model to assess their potential impact. This pipeline streamlines the process of AI-driven video generation and evaluation, enabling automatic popular micro-video content creation.
2.3 Prompt Enhancement
We demonstrate the Prompt Enhancement process in Figure 2, which enhances the input prompt for LLM inspired by human content creators.
Formally, let $x\in\mathcal{X}$ be the user prompt (e.g., "Create a video showcasing Goku’s most iconic transformations in Dragon Ball."). An embedding function $f_{\text{embed}}:\mathcal{X}\to\mathbb{R}^{d}$ maps $x$ to its vector representation:
$\mathbf{x}=f_{\text{embed}}(x).$
Given a predefined set of topic tags $\mathcal{T}$ (e.g., Food, Anime, Movie) with corresponding embeddings, a matching algorithm based on cosine similarity is used to select the most relevant tags:
$\mathcal{T}_{x}=\operatorname{arg\,max}_{t\in\mathcal{T}}sim\bigl{(}\mathbf{x},f_{\text{embed}}(t)\bigr{)},$
where $sim(\cdot,\cdot)$ denotes the cosine similarity. For each selected tag $t\in\mathcal{T}_{x}$, the system retrieves Top-K associated video samples from the test set of the Microlens dataset [^2] Since MMRA is trained on the training set of the Microlens dataset, we exclusively use the test set to prevent data leakage. and partitions them into positive examples $S^{+}_{t}$ and negative examples $S^{-}_{t}$ based on popularity. Finally, an enhancement function constructs the new prompt:
$x^{\prime}=\operatorname{Enhance}\Bigl{(}x,\ \bigcup_{t\in\mathcal{T}_{x}}\{S^{+}_{t},\,S^{-}_{t}\}\Bigr{)},$
which is then used to guide the video generation. The refined examples of the enhanced prompt templates are presented in Appendix Figure 10.
3 Experiment Setup
3.1 Dataset
To enhance comprehensiveness, we designed two types of user prompts as inputs for the LLM: Concrete and Abstract. The user prompts are primarily derived from five key categories—Anime, Delicacy, Daily Sharing, Film & Television, and Gaming. We reference the Microlens dataset Ni et al. (2023), as it is the largest available micro-video dataset. As detailed in Table 1, concrete prompts include specific elements that provide clear details, while abstract prompts describe videos in a more general and high-level manner. Specifically, we utilized ChatGPT-4o [^3] https://chatgpt.com/ to generate two corresponding user prompt datasets, each containing one hundred prompts [^4] Due to the high cost of video generation, overall two hundred prompts are similar size to existing video generation datasets.Jiang et al. (2023); Chivileva et al. (2024). These datasets are further divided into the aforementioned five categories, with each category comprising twenty prompts. This approach ensures a balanced and diverse set of prompts across all categories, catering to both specific and broad input styles for the LLM.
3.2 Evaluation
In this study, we evaluate the popularity of the generated videos using the state-of-the-art micro-video popularity prediction model, MMRA Zhong et al. (2024). Following Zhong et al. (2024), we define popularity as the number of comments. For robustness, we primarily use the median popularity score as the key metric. Following Zhong et al. (2024), we define popularity as the number of comments and use the median for robustness. This metric is greater than 0, with higher values indicating that the model generates prompts leading to greater median engagement, suggesting a higher potential for videos to achieve high engagement on the platform. In model-wise comparison sections such as Section 5 and Section 6, we primarily use Win-Rate to compare the performance of different models in predicting popularity. Specifically, given a user prompt, we generate micro-videos using both Model A and Model B and assess which model achieves higher popularity based on evaluation models. This pairwise comparison approach is widely adopted in the existing literature Zheng et al. (2024); Gao et al. (2024); Bianchi et al. (2024). More experiment details are in Apendix A.1.
4 LLM Pipeline for Popular Micro-videos Generation (RQ1)
To ensure a comprehensive evaluation, we consider three large language models (LLMs)—Llama-3.3-70B Touvron et al. (2023), ChatGPT-4o Hurst et al. (2024), and DeepSeek-V3 Liu et al. (2024a)—while setting CogVideoX Yang et al. (2024b) as the default video generation model. We utilize both concrete and abstract user prompt datasets as described in Section 3.1. Additionally, to better contextualize the popularity values of videos generated by the LLMPopcorn pipeline, we compare them with human-created videos from the Microlens dataset Ni et al. (2023). Since directly producing videos based on prompts would be prohibitively expensive, we approximate this comparison by matching each user prompt to the most semantically similar human-created video title using an embedding model (all-MiniLM-L12-v2). This approach mitigates the influence of the prompt’s thematic bias on video popularity, allowing for a closer examination of how the popularity of LLMPopcorn-generated videos compares to real human-created content.
As demonstrated in Table 2, compared to the Microlens human-created dataset, which achieves median popularity scores of 0.44 for concrete prompts and 0.71 for abstract prompts, the AI-generated content by LLMPopcorn demonstrates competitive performance, with certain models even surpassing human benchmarks using DeepSeek-V3. Among the evaluated models, DeepSeek-V3 stands out as the most robust and stable performer, achieving the highest median scores of 0.56 for concrete prompts and 0.73 for abstract prompts, highlighting its capability to generate content with strong engagement potential. However, while these models achieve performance on par with human-created videos on average, their absolute values remain relatively low. We hypothesize that this may stem from the limitations of current video generation models
@@ 4 LLM Pipeline for Popular Micro-videos Generation (RQ1)
Table 2: Comparison of popularity medians for video generation tasks (using CogVideoX-5b as the video generator). Human-created videos are included for reference. The popularity predictor provides predicted values of Human Crafted video for comparison. Concrete-M and Abstract-M are the most closely matched human videos to the same user prompt in the concrete and abstract datasets, respectively.
, which struggle to fully capture and convey the complexity of the given descriptions Liu et al. (2024b).
Category Abstract Concrete
Llama DeepSeek ChatGPT Llama DeepSeek ChatGPT
Anime 0.54 0.92 0.63 0.37 0.45 0.39
Daily Sharing 0.59 0.36 0.23 0.28 0.69 0.61
Delicacy 0.48 1.12 0.64 0.66 0.59 1.24
Film and Television Works 0.68 1.02 0.82 0.46 0.66 0.35
Game 0.38 0.49 0.55 0.35 0.30 0.56
Figure 3: Predicted popularity distributions for different LLMs. Evaluated using three LLMs: Llama-3.3-70B, DeepSeek-V3, and ChatGPT-4o. Concrete and Abstract represent the types of user prompt datasets.
The predicted popularity distributions are presented in Figure 3. For abstract prompts, the distributions of Llama-3.3-70B and ChatGPT-4o are highly concentrated near a popularity score of 0, with occasional outliers achieving higher scores. These rare successes highlight the overall instability of these models. In contrast, DeepSeek-V3 exhibits a smoother and more evenly distributed spread, with a higher proportion of videos falling into mid-to-high popularity ranges. However, in concrete datasets, the differences among the three models become negligible. This finding is further supported by DeepSeek-V3’s highest median score in abstract datasets, while its median score in concrete prompts remains comparable to the others.
Breaking down the analysis by category (Table 3), DeepSeek-V3 achieves the highest predicted popularity scores across most categories in the abstract dataset. Notably, it significantly outperforms other models in the Anime and Film and Television Works categories, while also demonstrating strong performance in Delicacy. However, its dominance is less pronounced in the Daily Sharing and Game categories, where its advantage is more limited.
@@
Table 4: Comparison of popularity win rates across different LLMs.
Video Generators Task Win Rate (%)
LTX Concrete 60
Abstract 61
Hunyuan Concrete 48
Abstract 52
Table 5: Comparison of popularity win rates across different video generation models.
(Answer to RQ1) LLMs facilitate micro-video creation by structuring descriptions and guiding Video generation models, with their effectiveness assessed using a pre-trained offline popularity predictor. While the absolute values remain modest, the median popularity of videos generated by state-of-the-art LLMs matches or even exceeds that of average human-created videos in our evaluation. It suggests that the proposed method has considerable potential for popular short video generation. These findings highlight promising directions for enhancing AI-generated video engagement.
5 Prompt-based Popularity Enhancement (RQ2)
The comparative results for Prompt Enhancement (PE) versus the basic LLMPopcorn prompt are summarized in Tables 4 and 5, using win rates as the evaluation metric. All win rates presented are relative values, derived from pairwise comparisons between PE and the basic prompt across different models and scenarios. Since these values are model-specific and do not represent absolute performance, they should not be used for direct cross-model comparisons.
Table 4 evaluates the PE’s effectiveness across three LLMs for both concrete and abstract video creation tasks. Notably, with Llama-3.3-70B, PE achieves a 56% win rate for concrete tasks and 52% for abstract tasks, surpassing basic LLMPopcorn. DeepSeek-V3 delivers competitive gains (55% on concrete tasks), but its relatively low abstract-task win rate (49%). In contrast, ChatGPT-4o exhibits balanced improvements, achieving a 53% win rate on abstract tasks, indicating strong compatibility with PE’s human-like creative process.
Table 5 further validates the impact of PE on two more video generation models. LTX exhibits the most significant improvements, achieving win rates of 60% on concrete tasks and 61% on abstract tasks. Hunyuan shows a slight performance gain on the Abstract dataset but performs marginally worse on the Concrete dataset. Overall, these results indicate that PE effectively enhances LLM-generated prompts for popular video generation models, despite some fluctuations.
LLM Model Comparison Concrete Dataset (%) Abstract Dataset (%)
BP WR PE WR BP WR PE WR

### v3

# LLMPopcorn: Exploring LLMs as Assistants for Popular Micro-video Generation

@@ Abstract
In an era where micro-videos dominate platforms like TikTok and YouTube, AI-generated content is nearing cinematic quality. The next frontier is using large language models (LLMs) to autonomously create viral micro-videos, a largely untapped potential that could shape the future of AI-driven content creation. To address this gap, this paper presents the first exploration of LLM-assisted popular micro-video generation (LLMPopcorn [^1] We selected popcorn as the icon for this paper because it symbolizes leisure and entertainment. This aligns with this study on leveraging LLMs as assistants for generating popular micro-videos, which are often consumed during leisure time.). Specifically, we empirically study the following research questions: (i) How can LLMs be effectively utilized to assist popular micro-video generation? (ii) To what extent can prompt-based enhancements optimize the LLM-generated content for higher popularity? (iii) How well do various LLMs and video generators perform in the popular micro-video generation task? Exploring these questions, we show that advanced LLMs like DeepSeek-V3 can generate micro-videos with popularity rivaling human content. Prompt enhancement further boosts results, while benchmarking highlights DeepSeek-V3 and R1 for LLMs, and LTX-Video and HunyuanVideo for video generation. This work advances AI-assisted micro-video creation and opens new research directions. The code is publicly available at https://github.com/GAIR-Lab/LLMPopcorn.
Index Terms—  Micro-video Generation, Video Popularity, LLM, LLMPopcorn
@@ 1 Introduction
Micro-videos (or short videos) have emerged as a crucial element of the digital economy, representing a multi-billion-dollar industry [10]. They have become an integral part of daily life for people worldwide, providing substantial commercial value for social media platforms and content creators. Popular content creators can receive significant revenue through their content [21], which underscores the ever-growing influence of micro-videos in modern society.
@@
Despite their widespread popularity and financial impact, producing popular micro-video content remains a costly and labor-intensive process. Professional filming, scripting, and editing require significant time and resources, which not all creators or businesses can afford. Driven by these challenges, we explore simplifying the micro-video creation process using the state-of-the-art AI solutions, with a particular focus on generating popular micro-videos.
@@
On one hand, the rapid advancement of Large Language Models (LLMs) like ChatGPT has unlocked new possibilities for content generation, demonstrating strong capabilities in tasks such as document summarization [14], programming [1], and recommendation [4, 7, 6, 28]. On the other hand, video generation models powered by diffusion techniques and neural rendering are transforming creative content production, enabling high-quality synthesis for applications in film production and interactive media [18]. Significant breakthroughs in LLMs and video generation models have prompted many studies [30, 12] to integrate LLMs to facilitate the capability of automatic video generation. For instance, VideoLLMs like VideoPoet [15] and GPT4Video [23] leverage LLMs and video generation models for multimodal video understanding and generation.
Current studies on LLMs for video generation overlook their role in creating popular micro-videos, focusing instead on quality metrics like resolution and duration. Recent research, including models like those in [19] and datasets like Microlens [20], emphasizes micro-video popularity prediction. However, the use of mainstream LLMs to generate highly popular micro-videos, leveraging their text generation and planning capabilities [25, 5], remains underexplored.
@@
Driven by the above insights, we present the first exploration of the research question of whether LLMs can assist in the popular micro-video creation. To address this research question systematically, we propose and investigate the following three Research Questions (RQs):
RQ1: How can LLMs be effectively utilized to assist popular micro-video generation?
@@
We present LLMPopcorn, a micro-video creation pipeline that uses LLMs to generate prompts for popular video production. To assess popularity, we integrate a state-of-the-art offline evaluator and compare three mainstream LLMs against human-created videos.
RQ2: To what extent can prompt-based enhancements optimize the content generated by LLMs for achieving greater popularity?
We propose a Prompt Enhancement combining Retrieval-Augmented Generation (RAG) [17] and chain-of-thought (CoT) prompting [24], inspired by human content creation.
@@
RQ3: How do various LLMs and video generators perform on popular micro-video generation?
We compare five LLMs and three video generation models using the LLMPopcorn pipeline.
@@
2 The Proposed LLMPopcorn
Fig. 1: Overview of the LLMPopcorn pipeline. An LLM processes a user query, generating a video title and prompt for a video generator to create a micro-video. The video and title are then evaluated for predicted popularity.
2.1 Problem Formulation
We consider a problem in which a user provides an input prompt $x\in\mathcal{X}$. An LLM processes this prompt and generates both a video title and a video prompt, which we denote by $(t,p)=f(x)$, where $t$ belongs to the set of titles $\mathcal{T}$ and $p$ belongs to the set of video prompts $\mathcal{P}$. Next, a pre-trained video generation model $g$ uses the generated prompt to produce an AI-created micro-video $v\in\mathcal{V}$, that is, $v=g(p)$. An offline video popularity predictor $h$ then assigns a popularity score $s\in\mathbb{R}$ to the generated video, so that $s=h(v,t)$. Overall, the entire pipeline can be represented as the composite function: $F(x)=h\bigl(g(f(x))\bigr).$
The objective is to optimize the LLM function $f$ (with $g$ and $h$ fixed) to maximize the expected popularity score of the generated videos. Formally, we wish to achieve: $\max_{f\in\mathcal{F}}\;\mathbb{E}_{x\sim\mathcal{X}}\left[h\bigl(g(f(x))\bigr)\right],$ where $\mathcal{F}$ denotes the set of all possible configurations of the LLM. This formulation captures the aim of leveraging LLMs to enhance the creation of popular AI-generated micro-videos.
2.2 Pipeline
Based on the problem formulation, we propose a new LLMPopcorn pipeline in Figure 1 that begins by constructing user prompts, categorized into concrete and abstract prompts, which are constructed by rules into inputs for an LLM. The LLM serves as a video generation assistant, producing video titles and video generation prompts based on these inputs. These video generation prompts are then fed into a video generator to create corresponding videos. The generated videos are combined with their titles to form final short videos, which are subsequently evaluated using a pre-trained video popularity prediction model to assess their potential impact. This pipeline streamlines the process of AI-driven video generation and evaluation, enabling automatic popular micro-video content creation.
Fig. 2: An overview of the Prompt Enhancement (PE) process. PE enables the LLM reviewing relevant micro-videos from the database and engaging in chain-of-thought reasoning.
2.3 Prompt Enhancement
We demonstrate the Prompt Enhancement process in Figure 2, which enhances the input prompt for LLM inspired by human content creators. Formally, let $x\in\mathcal{X}$ be the user prompt (e.g., “Create a video showcasing Goku’s most iconic transformations in Dragon Ball.”). An embedding function $f_{\text{embed}}:\mathcal{X}\to\mathbb{R}^{d}$ maps $x$ to its vector representation: $\mathbf{x}=f_{\text{embed}}(x).$ Given a predefined set of topic tags $\mathcal{T}$ (e.g., Food, Anime, Movie) with corresponding embeddings, a matching algorithm based on cosine similarity is used to select the most relevant tags: $\mathcal{T}_{x}=\operatorname{arg\,max}_{t\in\mathcal{T}}sim\bigl(\mathbf{x},f_{\text{embed}}(t)\bigr),$ where $sim(\cdot,\cdot)$ denotes the cosine similarity. For each selected tag $t\in\mathcal{T}_{x}$, the system retrieves Top-K associated video samples from the test set of the Microlens dataset [^2] Since MMRA is trained on the training set of the Microlens dataset, we exclusively use the test set to prevent data leakage. and partitions them into positive examples $S^{+}_{t}$ and negative examples $S^{-}_{t}$ based on popularity. Finally, an enhancement function constructs the new prompt: $x^{\prime}=\operatorname{Enhance}\Bigl(x,\ \bigcup_{t\in\mathcal{T}_{x}}\{S^{+}_{t},\,S^{-}_{t}\}\Bigr),$ which is then used to guide the video generation.
3 Experiment Setup
Dataset. To enhance comprehensiveness, we designed two types of user prompts as inputs for the LLM: Concrete and Abstract. The user prompts are primarily derived from five key categories—Anime, Delicacy, Daily Sharing, Film & Television, and Gaming. We reference the Microlens dataset [20], as it is the largest available micro-video dataset. Concrete prompts include specific elements that provide clear details, while abstract prompts describe videos in a more general and high-level manner. Specifically, we utilized ChatGPT-4o [13] to generate two corresponding user prompt datasets, each containing one hundred prompts, based on instructions to include either concrete elements or general content. [^3] Due to the high cost of video generation, overall two hundred prompts are similar size to existing video generation datasets.[3] These datasets are further divided into the aforementioned five categories, with each category comprising twenty prompts. This approach ensures a balanced and diverse set of prompts across all categories, catering to both specific and broad input styles for the LLM.
Evaluation. In this study, we assess the popularity of generated micro-videos using the MMRA model [29], defining popularity as the median number of comments for robustness, as per [29]. Higher values indicate greater median engagement, suggesting videos with higher engagement potential. In model-wise comparisons (e.g., Section 5 and Section 6), we use Win-Rate [8] to evaluate model performance in predicting popularity, comparing micro-videos generated by Model A and Model B for a given user prompt. This pairwise approach is common in the literature [8]. Additionally, we conducted a user study following [2], involving five expert volunteers with master’s degrees and over five years of experience in micro-video consumption and creation.
@@ 3 Experiment Setup
Llama DeepSeek GPT Llama DeepSeek GPT
Anime 0.54 0.92 0.63 0.37 0.45 0.39
Daily Sharing 0.59 0.36 0.23 0.28 0.69 0.61
Delicacy 0.48 1.12 0.64 0.66 0.59 1.24
Film 0.68 1.02 0.82 0.46 0.66 0.35
Game 0.38 0.49 0.55 0.35 0.30 0.56
Table 2: Comparison of popularity medians for video generation tasks (using CogVideoX-5b as the video generator). Human-created videos are included for reference. The popularity predictor provides predicted values of Human Crafted video for comparison. Concrete-M and Abstract-M represent the human videos from the Microlens dataset that are most closely matched to the same user prompt in the concrete and abstract datasets, respectively.
@@
Fig. 3: Predicted popularity distributions for different LLMs. Y-axis: frequency. X-axis: predicted popularity.
Table 3: Comparison of popularity for Prompt Enhancement’s win rates across different LLMs.
@@
Table 4: User study for Prompt Enhancement vs. Basic LLMPopcorn prompt.
LLM Task Win Rate (%)
User Evaluation Concrete 66.67
Abstract 62.50
Implementation details In this paper, we utilized five state-of-the-art LLMs, including two open-source models, Llama-3.3-70B and Qwen-2.5-72B, and three API-based models, DeepSeek-V3, DeepSeek-R1, and ChatGPT-4o. For efficiency, we apply 4-bit quantization to the open-source LLMs. Additionally, we incorporated three widely-used open-source video generation models: CogVideoX-5B, LTX-Video, and HunyuanVideo. All model weights were obtained from the HuggingFace platform [^4] https://huggingface.co/ to ensure consistency and reproducibility. Notably, all open-source LLMs used in this study were instruction-tuned versions. We ensured reproducibility by fixing the random seed for opensourced models, allowing for consistent results. We adopt Faiss [^5] https://github.com/facebookresearch/faiss as a library for matching and indexing. For the prompt enhancement related to RQ2, we search the retrieval-augmented generation (RAG) size from [10, 50, 100]. The video generation prompt contains existing video titles and video captions paired with their popularity. All experiments are conducted on one H100 80G GPU. Due to the space limit, the prompt templates and user query datasets used in this paper can be checked from the code repository.
4 Popular Micro-videos Generation (RQ1)
LLM Model Comparison Concrete Abstract
 BP PE  BP PE